\int \:\frac{1+sin\left(mx\right)}{cos^2\left(mx\right)}dx
1
Expert's answer
2020-10-05T10:14:28-0400

"\int \:\frac{1+sin\left(mx\right)}{cos^2\left(mx\right)}dx"

replacement

"u = mx; \frac{du}{dx}=m;" "dx = \frac{1}{m}du;"


"\frac{1}{m}\int\frac{sin(u)+1}{cos^2(u)}du =\frac{1}{m} \int\frac{sin(u)}{cos^2(u)}du+" "\frac{1}{m}\int\frac{1}{cos^2(u)}du" =  "=\frac{1}{m} \int\frac{sin(u)}{cos^2(u)}du+\frac{1}{m}tan(u)"


"\frac{1}{m}\int\frac{sin(u)}{cos^2(u)}du" ; replacement "v=cosu;\frac{dv}{du}=-sin(u);du=-\frac{1}{sin(u)}dv"

"\frac{1}{m}\int\frac{sin(u)}{cos^2(u)}du =-\frac{1}{m}\int\frac{1}{v^2}dv=-\frac{1}{mv}" ="-\frac{1}{mcos(u)}"


"\frac{1}{m}\int\frac{sin(u)+1}{cos^2(u)}du =\frac{1}{m}tan(u)-- \frac{1}{mcos(u)}"

reverse replacement u = mx

"\int \:\frac{1+sin\left(mx\right)}{cos^2\left(mx\right)}dx =" "\dfrac{\tan\left(mx\right)}{m}+\dfrac{1}{m\cos\left(mx\right)} +C"


Answer: "\dfrac{\tan\left(mx\right)}{m}+\dfrac{1}{m\cos\left(mx\right)} +C"